### v5

@@ 5 L_List,L0_List,C_List,Tracked_Alignment=SRG_Function(L,L0,q,IterNum)
1G=nx.random_graphs.barabasi_albert_graph(1000,4) # Generate a a random BA network with 1000 units
@@
4L_List,L0_List,C_List,TrackedNodeAlignment=SRGFlow(G,q=1,p=1,L_Type=’MOL’,IterNum=5) # Run a SRG for 5 iterations, which renormalize the system on the 1-order based on the 1-order interactions
@@
6L_List,L0_List,C_List,TrackedNodeAlignment=SRGFlow(G,q=2,p=1,L_Type=’MOL’,IterNum=5) # Run a SRG for 5 iterations, which renormalize the system on the 1-order based on the 1-order interactions
@@
8L_List,L0_List,C_List,TrackedNodeAlignment=SRGFlow(G,q=3,p=1,L_Type=’MOL’,IterNum=5) # Run a SRG for 5 iterations, which renormalize the system on the 1-order based on the 1-order interactions
@@
11L_List,L0_List,C_List,TrackedNodeAlignment=SRGFlow(G,q=1,p=1,L_Type=’MOL’,IterNum=5) # Run a SRG for 5 iterations, which renormalize the system on the 1-order based on the 1-order interactions
@@
13L_List,L0_List,C_List,TrackedNodeAlignment=SRGFlow(G,q=2,p=1,L_Type=’MOL’,IterNum=5) # Run a SRG for 5 iterations, which renormalize the system on the 1-order based on the 1-order interactions
@@
15L_List,L0_List,C_List,TrackedNodeAlignment=SRGFlow(G,q=3,p=1,L_Type=’MOL’,IterNum=5) # Run a SRG for 5 iterations, which renormalize the system on the 1-order based on the 1-order interactions
@@ 61 HO_A[i,j] += 1
63 HO_D = sum(HO_A)
@@
69 HO_A[i,j] += 1
70 HO_A += HO_A.T
71 HO_D = sum(HO_A)
72 L = np.diag(HO_D) - HO_A
@@ 93 L=L[1:,1:]
95
96def SRG_Flow(G,q,p,L_Type,IterNum):
@@
98 L=HighOrderLaplician(A, L_Type, Order=q)
99 L0=HighOrderLaplician(A, L_Type, Order=p)
100 L_List,L0_List,C_List,Tracked_Alignment=SRG_Function(L,L0,q,IterNum)
101 return L_List,L0_List,C_List,Tracked_Alignment
102
103
104def SRG_Function(L,L0,q,IterNum):
@@ 99 L0=HighOrderLaplician(A, L_Type, Order=p)
106 # L is the initial guiding high order Laplician representation, which can be the Multiorder Laplacian operator or the high-order path Laplacian
107 # L0 is the initial guided high order Laplician representation, which can be the Multiorder Laplacian operator or the high-order path Laplacian
108
@@
110 # L_List is the list of the guiding q-order Laplacian over renormalization steps
111 # L0_List is the list of the guided p-order Laplacian over renormalization steps
@@
115 Init_Sub_Ls,_=NetworkInfo(L)
116 L_List=[L]
117 L0_List=[L0]
118 tau_Vec=np.zeros(len(Init_Sub_Ls))
@@
120 for ID in range(len(Init_Sub_Ls)):
121 if np.size(Init_Sub_Ls[ID],0)>1:
122 tau,CVector=TauSelection(Init_Sub_Ls[ID])
@@
128 Current_L=L_List[Iter-1]
129 Current_L0=L0_List[Iter-1]
130 Current_G0=nx.from_numpy_array(np.abs(Current_L0)-np.diag(np.diag(Current_L0)))
131 Sub_Ls, Sub_NodeIDs=NetworkInfo(Current_L)
132 NewSub_Ls=[]
@@
134 for SLID in range(len(Sub_Ls)):
135 SL=Sub_Ls[SLID]
136 SNodeID = Sub_NodeIDs[SLID]
137 if np.size(SL,0)==1:
@@
139 NewSub_Ls.append(SL)
@@
143 Evals, _ = np.linalg.eig(SL)
144 tau=tau_Vec[SLID]
145 Eval_idx = np.where(np.real(Evals)>1/tau)[0]
146
147 Rho = expm(-tau*SL)/np.trace(expm(-tau*SL))
148 rho_Evals, rho_Evecs = np.linalg.eig(Rho)
149 rho = np.zeros((Rho.shape[0],Rho.shape[1]))
150 for ID in Eval_idx:
151 Evec = rho_Evecs[:,ID].reshape(-1,1)
152 rho += np.real(rho_Evals[ID]*np.matmul(Evec,Evec.T))
153 adj1 = np.zeros((rho.shape[0],rho.shape[1]))
154 for i in range(SL.shape[0]):
155 for j in range(SL.shape[1]):
156 if i==j:
157 continue
158 elif ((rho[i][j]>=rho[j][j]) or (rho[i][j]>=rho[i][i])):
159 adj1[i][j]=1
160 adj1[j][i]=1
161 RefG = nx.from_numpy_array(adj1) # reference graph
162 # delete false edges not in subgraph of SL
163 Current_SG=nx.from_numpy_array(np.abs(SL)-np.diag(np.diag(SL)))
164 Potential_Clusters=[list(c) for c in list(nx.connected_components(RefG))]
165 Edge_To_Remove=[]
166 for ID1 in range(len(Potential_Clusters)):
167 Unit_list=Potential_Clusters[ID1]
168 if len(Unit_list)>1:
169 H = nx.induced_subgraph(Current_SG,Unit_list)
170 Potential_H = nx.induced_subgraph(RefG,Unit_list)
171 Wrong_Edge=list(set(list(Potential_H.edges))-set(list(H.edges)))
172 Edge_To_Remove.extend(Wrong_Edge)
173 for Wrong_Edge in Edge_To_Remove:
174 RefG.remove_edge(*Wrong_Edge)
@@
176 # contracting SG nodes
177 for Nodes in Clusters:
178 Node1 = Nodes[0]
179 for node in Nodes[1:]:
180 Current_SG = nx.contracted_nodes(Current_SG, Node1, node, self_loops=False)
181 mapping = {old_label:new_label for new_label, old_label in enumerate(Current_SG.nodes())}
182 Current_SG = nx.relabel_nodes(Current_SG, mapping)
183 New_SA = nx.adjacency_matrix(Current_SG).toarray()
184 # contracting G0 nodes
@@
189 for node in SNodes[1:]:
190 Current_G0 = nx.contracted_nodes(Current_G0, Node1, node, self_loops=False)
191
192 New_SL = np.diag(np.sum(New_SA,axis=0)) - New_SA
193 NewSub_Ls.append(New_SL)
@@
196
197 mapping = {old_label:new_label for new_label, old_label in enumerate(Current_G0.nodes())}
198 Current_G0 = nx.relabel_nodes(Current_G0, mapping)
199 New_A0 = nx.adjacency_matrix(Current_G0).toarray()
200 New_L0 = np.diag(np.sum(New_A0,axis=0)) - New_A0
201 New_L=NetworkUnion(NewSub_Ls)
202 print(’Nn:’, New_L.shape[0])
203
204 L_List.append(New_L)
205 L0_List.append(New_L0)
@@
207 Tracked_Alignment=TrackingUnitID(All_Alignment,L.shape[0])
208 return L_List,L0_List,C_List,Tracked_Alignment
209
210def TauSelection(L):
@@
216 TauVec=np.logspace(-2, 2, 200) # reconmennd tau in (1e-2,1e0) for q>1
@@
220 MatrixExp=spy.linalg.expm(-Poss_tau*L)
@@
238
239def TrackingUnitID(All_Alignment,UnitNum):

### v7

@@ 5 Lq_List,Lp_List,Gq_List,Gp_List,C_List,Tracked_Alignment=SRG_Function(Lq,Lp,q,IterNum)
1G=nx.random_graphs.barabasi_albert_graph(1000,3) # Generate a random BA network with 1000 units
@@
4Lq_List,Lp_List,Gq_List,Gp_List,C_List,Tracked_Alignment=SRG_Flow(G,q=1,p=1,L_Type=’MOL’,IterNum=5) # Run a SRG for 5 iterations, which renormalize the system on the 1-order based on the 1-order interactions
@@
6Lq_List,Lp_List,Gq_List,Gp_List,C_List,Tracked_Alignment=SRG_Flow(G,q=2,p=1,L_Type=’MOL’,IterNum=5) # Run a SRG for 5 iterations, which renormalize the system on the 1-order based on the 2-order interactions
@@
8Lq_List,Lp_List,Gq_List,Gp_List,C_List,Tracked_Alignment=SRG_Flow(G,q=3,p=1,L_Type=’MOL’,IterNum=5) # Run a SRG for 5 iterations, which renormalize the system on the 1-order based on the 3-order interactions
@@
11Lq_List,Lp_List,Gq_List,Gp_List,C_List,Tracked_Alignment=SRG_Flow(G,q=1,p=1,L_Type=’HOPL’,IterNum=5) # Run a SRG for 5 iterations, which renormalize the system on the 1-order based on the 1-order interactions
@@
13Lq_List,Lp_List,Gq_List,Gp_List,C_List,Tracked_Alignment=SRG_Flow(G,q=2,p=1,L_Type=’HOPL’,IterNum=5) # Run a SRG for 5 iterations, which renormalize the system on the 1-order based on the 2-order interactions
@@
15Lq_List,Lp_List,Gq_List,Gp_List,C_List,Tracked_Alignment=SRG_Flow(G,q=3,p=1,L_Type=’HOPL’,IterNum=5) # Run a SRG for 5 iterations, which renormalize the system on the 1-order based on the 3-order interactions
@@ 61 HO_A[i,j] += 1
63 HO_D = np.zeros(Num)
64 for Simplex in Cliques:
65 for [i] in itertools.combinations(Simplex,1):
66 HO_D[i] += 1
@@
72 HO_B[i,j] += np.math.factorial(Order-1)
73 HO_B += HO_B.T
74 HO_P = sum(HO_B)
75 L = (np.diag(HO_P) - HO_B)/Order
@@ 96 L=L[1:,1:]
98
99def SRG_Flow(G,q,p,L_Type,IterNum):
@@
101 Lq=HighOrderLaplician(A, L_Type, Order=q)
102 Lp=HighOrderLaplician(A, L_Type, Order=p)
103 Lq_List,Lp_List,Gq_List,Gp_List,C_List,Tracked_Alignment=SRG_Function(Lq,Lp,q,IterNum)
104 return Lq_List,Lp_List,Gq_List,Gp_List,C_List,Tracked_Alignment
105
106
107def SRG_Function(Lq,Lp,q,IterNum):
@@
109 # Lq is the initial guiding high order Laplician representation, which can be the Multiorder Laplacian operator or the high-order path Laplacian
110 # Lp is the initial guided high order Laplician representation, which can be the Multiorder Laplacian operator or the high-order path Laplacian
111
@@
114 # Lq_List is the list of the guiding q-order Laplacian over renormalization steps
115 # Lp_List is the list of the guided p-order Laplacian over renormalization steps
116 # Gq_List is the list of the guiding q-order network sketches over renormalization steps
117 # Gp_List is the list of the guided p-order network sketches over renormalization steps
@@
121 Gq=nx.from_numpy_array(np.abs(Lq)-np.diag(np.diag(Lq)))
122 Gp=nx.from_numpy_array(np.abs(Lp)-np.diag(np.diag(Lp)))
123 Lq_List=[Lq]
124 Lp_List=[Lp]
125 Gq_List=[Gq]
126 Gp_List=[Gp]
@@
128 Init_Sub_Lqs,_=NetworkInfo(Lq)
129 tau_Vec=np.zeros(len(Init_Sub_Lqs))
130 for ID in range(len(Init_Sub_Lqs)):
131 if np.size(Init_Sub_Lqs[ID],0)>1:
132 tau,CVector=TauSelection(Init_Sub_Lqs[ID])
@@
138 ## step 1: initiate Laplacian operators and associated high-order network sketches in the k-th iteration
139 Origin_Lq,Origin_Lp=Lq_List[Iter-1],Lp_List[Iter-1]
140 Origin_Gq,Origin_Gp=Gq_List[Iter-1],Gp_List[Iter-1]
141 Origin_Ap=nx.adjacency_matrix(Origin_Gp).toarray()
142 New_Ap=copy.deepcopy(Origin_Ap)
143 New_Lp=copy.deepcopy(Origin_Lp)
144 Sub_Lqs, Sub_NodeIDs=NetworkInfo(Origin_Lq)
145 Sub_Gqs=[Origin_Gq.subgraph(SNodeID).copy() for SNodeID in Sub_NodeIDs]
146 NewSub_Lqs=[]
147 NewSub_Gqs=[]
148 New_Snodes_list=[]
@@
150 for SID in range(len(Sub_Lqs)):
151 SLq=Sub_Lqs[SID]
152 SGq=Sub_Gqs[SID]
153 SNodeID = Sub_NodeIDs[SID]
154 if np.size(SLq,0)==1:
@@
156 NewSub_Lqs.append(SLq)
157 NewSub_Gqs.append(SGq)
158 New_Snodes_list.extend(SNodeID)
@@
162 ## step 2: search targets to coarse grain in the real space
163 Rho = expm(-tau*SLq)/np.trace(expm(-tau*SLq))
164 diagonal_min = np.minimum(Rho.diagonal().reshape(-1, 1), Rho.diagonal())
165 Rho_prime=Rho/diagonal_min
166 Rho_prime=(Rho_prime+Rho_prime.T)/2
167 Ref_adj=(Rho_prime>1).astype(’int’)
168 # delete false edges not in subgraph of SLq
169 Origin_SAq=nx.adjacency_matrix(SGq).toarray()
170 Ref_mask=(Origin_SAq!=0).astype(int)
171 Ref_adj=Ref_adj*Ref_mask
172 RefG = nx.from_numpy_array(Ref_adj) # reference graph
@@
174 n_k=len(Clusters)
175 ## step 3: implement a renormalization procedure for q-order and p-order network sketches in the real space
176 # contracting SGq nodes
177 New_SAq=copy.deepcopy(Origin_SAq)
178 New_nodes=[Nodes[0] for Nodes in Clusters]
179 for i in range(n_k):
180 iNodes=Clusters[i]
181 for j in range(i+1,n_k):
182 jNodes=Clusters[j]
183 new_weight=sum(Origin_SAq[iNodes,:][:,jNodes].reshape(-1))
184 New_SAq[iNodes[0],jNodes[0]]=new_weight
185 New_SAq[jNodes[0],iNodes[0]]=new_weight
186 New_SAq=New_SAq[New_nodes,:][:,New_nodes]
187 New_SGq=nx.from_numpy_array(New_SAq)
188 NewSub_Gqs.append(New_SGq)
189 # contracting Gp nodes
190 New_Snodes=[SNodeID[Nodes[0]] for Nodes in Clusters]
191 New_Snodes_list.extend(New_Snodes)
192 Non_Nodes=[x for x in list(range(New_Ap.shape[0])) if x not in SNodeID]
193 for i in range(n_k):
194 iSNodes = [SNodeID[node] for node in Clusters[i]]
195 New_Ap[iSNodes[0],Non_Nodes]=np.sum(Origin_Ap[iSNodes,:][:,Non_Nodes],axis=0) # merge edges from different connected components
196 for j in range(i+1,n_k):
197 jSNodes = [SNodeID[node] for node in Clusters[j]]
198 new_weight=sum(Origin_Ap[iSNodes,:][:,jSNodes].reshape(-1))
199 New_Ap[iSNodes[0],jSNodes[0]]=new_weight
200 New_Ap[jSNodes[0],iSNodes[0]]=new_weight
201 ## step 4: search modes to reduce in the moment space
202 Evalqs, Evecqs=np.linalg.eig(SLq)
203 tau=tau_Vec[SID]
204 Evalq_idx=np.where(np.real(Evalqs)<1/tau)[0]
205 m_k=len(Evalq_idx) if len(Evalq_idx)>=n_k else n_k
206 ## step 5: calculate the re-scaled contributions of long-range eigenvectors
207 SQq = np.zeros((SLq.shape[0],SLq.shape[1]))
208 Evalq_idx=Evalqs.argsort()[:m_k]
209 for ID in Evalq_idx:
210 Evecq = Evecqs[:,ID].reshape(-1,1)
211 SQq += np.real(Evalqs[ID]*np.matmul(Evecq,Evecq.T))
212
213 SLp=Origin_Lp[SNodeID,:][:,SNodeID]
214 Evalps, Evecps = np.linalg.eig(SLp)
215 Evalp_idx=Evalps.argsort()[:m_k]
216 SQp = np.zeros((SLp.shape[0],SLp.shape[1]))
217 for ID in Evalp_idx:
218 Evecp = Evecps[:,ID].reshape(-1,1)
219 SQp += np.real(Evalps[ID]*np.matmul(Evecp,Evecp.T))
220 ## step 6: coarse grain q-order and p-order Laplacian in the momentum space
221 # calculate new SLq and SLp
222 New_SLq = np.zeros((n_k,n_k))
223 New_SLp = np.zeros((n_k,n_k))
224 for i in range(n_k):
225 iNodes=Clusters[i]
226 iSNodes = [SNodeID[node] for node in Clusters[i]]
227 New_Lp[iSNodes[0],Non_Nodes]=np.sum(Origin_Lp[iSNodes,:][:,Non_Nodes],axis=0) # merge edges from different connected components
228 for j in range(i,n_k):
229 jNodes=Clusters[j]
230 new_weight=sum(SQq[iNodes,:][:,jNodes].reshape(-1))
231 new_weight0=sum(SQp[iNodes,:][:,jNodes].reshape(-1))
232 New_SLq[i,j]=New_SLq[j,i]=new_weight
233 New_SLp[i,j]=New_SLp[j,i]=new_weight0
234 # remove false edges in SLq and SLp
235 maskq=(New_SAq!=0).astype(int)
236 New_SLq=New_SLq*maskq
237 maskp=(New_Ap[New_Snodes,:][:,New_Snodes]).astype(int)
238 New_SLp=New_SLp*maskp
239 # ensure the sum of each row of q-order Laplacian operator is zero
240 for i in range(n_k):
241 tmp=list(range(n_k))
242 tmp.remove(i)
243 New_SLq[i,i]=-np.sum(New_SLq[i,tmp])
244 NewSub_Lqs.append(New_SLq)
245
246 for i in range(n_k):
247 iNode1=New_Snodes[i]
248 for j in range(i+1,n_k):
249 jNode1=New_Snodes[j]
250 New_Lp[iNode1,jNode1]=New_SLp[i,j]
251 New_Lp[jNode1,iNode1]=New_SLp[j,i]
252 # track node alignment
@@
259
260 New_Gq=nx.disjoint_union_all(NewSub_Gqs)
261 New_Ap=New_Ap[New_Snodes_list,:][:,New_Snodes_list]
262 New_Ap=np.maximum(New_Ap,New_Ap.T) # gaurantee the symmetry
263 New_Gp=nx.from_numpy_array(New_Ap)
264 New_Lp = New_Lp[New_Snodes_list,:][:,New_Snodes_list]
265 New_Lp=np.minimum(New_Lp,New_Lp.T) # gaurantee the symmetry
266 # ensure the sum of each row of p-order Laplacian operator is zero
267 for i in range(New_Lp.shape[0]):
268 tmp=list(range(New_Lp.shape[0]))
269 tmp.remove(i)
270 New_Lp[i,i]=-np.sum(New_Lp[i,tmp])
271 New_Lq=NetworkUnion(NewSub_Lqs)
272 print(New_Lq.shape)
273
274 Lq_List.append(New_Lq)
275 Lp_List.append(New_Lp)
276 Gq_List.append(New_Gq)
277 Gp_List.append(New_Gp)
@@
279 Tracked_Alignment=TrackingUnitID(All_Alignment,Lq.shape[0])
280 return C_List,Lq_List,Lp_List,Gq_List,Gp_List,All_Alignment
281
282def TauSelection(L):
@@
288 TauVec=np.logspace(-2, np.log10(50), 200)
289 A=np.abs(L)-np.diag(np.diag(L))
290 n_ks=[]
@@
294 MatrixExp=expm(-Poss_tau*L)
@@
297
298 diagonal_min = np.minimum(Rho.diagonal().reshape(-1, 1), Rho.diagonal())
299 Rho_prime=Rho/diagonal_min
300 Rho_prime=(Rho_prime+Rho_prime.T)/2
301 Ref_adj=(Rho_prime>1).astype(’int’)
302 # delete false edges not in subgraph of SL
303 Ref_mask=(A!=0).astype(int)
304 Ref_adj=Ref_adj*Ref_mask
305 RefG = nx.from_numpy_array(Ref_adj) # reference graph
306 n_ks.append(nx.number_connected_components(RefG))
@@
320
321 n_ks=np.array(n_ks)
322 n_ks=n_ks[np.where(~np.isnan(CVector))[0]]
323 dn_ks=np.abs(np.diff(n_ks)/np.diff(TauVec))
324 bi_tau=TauVec[np.argmax(dn_ks)]
325 tau=tau if tau>bi_tau else bi_tau
@@
328
329def TrackingUnitID(All_Alignment,UnitNum):